Accelerator Update for November 17, 2008 - November 19, 2008
Monday Nov 17
The day shift began with Operations monitoring the Tevatron (TeV) store 6566 (150.31E30), the Recycler stash (141.5E10), and the Antiproton Source stack (13.2E10), with the Linac down due to RF station (LRF5) problems, with the Booster and Main Injector (MI) waiting on Linac, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, with NuMI down due to a horn problem, and with the MiniBooNE experiment off due to Linac problems.
Operations prepared NuMI for an access at 9:31 AM to investigate its horn problem.
Operations reported that the LRF5 work was finished. Operations resumed sending beam to Pbar and MiniBooNE at 10:04 AM.
A MI expert began a 45-minute coalescing check at 3:10 PM.
Operations reported at 4 PM that the Recycler had stashed antiprotons five times during the day shift.
Operations reported at 8:10 PM that a TeV power supply (VB47) was having regulation problems. The orbit stabilization program corrected other power supplies to keep up when VB47 ramped to zero, but couldn’t when the supply ramped back up. This distorted the orbit and caused the quench. Operators changed out its control and regulation chassis.
Operations put the TeV through a dry squeeze at 11:05 PM.
Operations put the TeV through a wet squeeze at 11:43 PM.
Operations reported at 11:56 PM that the Recycler had stashed six times during the evening shift.
A Recycler expert reported at 11:59 PM that they found a BPM with an unstable readback feeding bad info into a compensation-loop. This problem was worked around as a temporary measure.
Tuesday Nov 18
The midnight shift began with Operations monitoring the stash (376.3E10), with the Pbar stacking (60.4E10), with NuMI off due to horn problems, and with MiniBooNE taking beam.
Operations put the TeV through a dry squeeze at 12:28 AM.
Operations put the TeV into shot setup at 12:52 AM.
Operations prepared MI for an access at 2:16 AM to investigate a Recycler kicker problem.
Operations established store 6569 at 2:21 AM with an initial luminosity of 306.6E30.
The MI access ended and Operations resumed sending beam to Pbar and MiniBooNE at 4:41 AM.
At 7:24 AM, Operations reported that a Linac Klystron RF station (KRF5) had tripped off twice. It reset.
Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.
At 10:04 AM, the TeV sector C4 air compressor tripped off. It reset. Operations notified experts who installed a portable compressor and brought it online while they investigated and made repairs.
At 2:37 PM, Operations reported that the TeV sector E4 air compressor’s pressure had dropped. Operations informed an expert.
At 3:30 PM, a TeV RF station (TRF3) tripped off and caused a loss of luminosity, about 6E30. Operations contacted an expert.
A NuMI expert reported at 3:55 PM that there was a problem with the current carrying stripline for horn #2. The horn might have to be replaced, which would take around 3 weeks. Their investigations continued.
Operations reported at 4:07 PM that the Recycler had stashed seven times during the day shift.
A TeV RF expert reported at 5:28 PM that he found nothing wrong with TRF3. He changed out the flow turbine and the magnetic pickup for the station just to be safe.
Operations terminated store 6569 at 6:12 PM.
Operations put the TeV through a dry squeeze at 7:12 PM.
Operations prepared MiniBooNE for an access at 7:58 PM to investigate the horn LCW pump.
Operations put the TeV through a wet squeeze at 8 PM.
At 8:12 PM, Operations reported errors associated with an orbit smoothing control card. Operators replaced the card.
At 9:06 PM, a MiniBooNE expert reported that their horn LCW pump had failed; it will have to be replaced. The experts began arranging for the work.
Operations put the TeV through a wet squeeze at 9:11 PM.
Operations put the TeV into shot setup at 9:34 PM.
At 10:22 PM, Operations reported TCLK errors on two TeV corrector magnet power supplies (VA12 and VF47). Operators had changed the control card and regulator chassis for both devices earlier in the evening. They discovered that they had used a bad card for VF47 during the last change. They replaced this card again.
Operations put the TeV through another dry squeeze at 11:25 PM.
Operations restarted shot setup at 11:53 PM.
Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.
Wednesday Nov 19
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (351.2E10), with Pbar stacking (117.1E10), with NuMI off due to horn problems, and with MiniBooNE off to replace a LCW pump for its horn.
MiniBooNE came out of access at 12:42 AM. They had replaced the horn LCW pump.
Operations resumed sending beam to MiniBooNE at 1:17 AM.
Operations established store 6574 at 1:27 AM with an initial luminosity of 305.51E30.
Operations reported at 7:57 AM that the Recycler had stashed six times during the midnight shift.
The Plans for Wednesday
The plans for today are to give CDF a four to six hour access to replace a crate power supply. During this access, the Run Coordinator will allow other accesses and maintenance work. She will also allow cryo system experts to overhaul the TeV sector C3 wet engine.
Comments and Suggestions
|last modified 11/20/2008 email Fermilab|
|Security, Privacy, Legal|